How mathematically prove that in meter bridge experiment accuracy is greater if the null point is obtain in middle portion of wire

Grade:7

2 Answers

Vikas TU
7 years ago
Reposting the answer as have been raised twice,
 
Null point should be at the middle of one third of the wire to obtain the greater accuracy.
Keeping in mind the sensitivity of the meter bridge it has to be done.
Beacuse at this point the ratio of the resistance sof the arms tends to nearly equal.
Sarasij Basu Mallick
7 years ago
p/q=(l/100-l) \Rightarrow dp/p=(((100-L)+l^2))/(100-l))\Rightarrow dp/p=(100-l+l^2)/(100l-l))\Rightarrow dp/p=-((100-l+l^2)/(l^2-100l+50^2-50^2))\Rightarrow dp/p= ((l-l^2-100)/((l-50)^2-50^2))
 
hence dp/p means error in measuring.we have to minimize it. so (l/50)^2=0 which implies l=50
 meter bridge is 100cm in length.so it is the most accuriate at l=50cm